OpenStudy (anonymous):

True or False: (1/2)(2n + 2) + -2 = n

OpenStudy (anonymous):

False

OpenStudy (anonymous):

cuzs theres no solution.

OpenStudy (anonymous):

(1/2)*(2n+2) - 2 = n (n+1)-2 = n n-1=n 0=1 False

OpenStudy (anonymous):

thank you!